### v1

# Bongard-RWR+: Real-World Representations of Fine-Grained Concepts in Bongard Problems

Abstract
Bongard Problems (BPs) provide a challenging testbed for abstract visual reasoning (AVR), requiring models to identify visual concepts from just a few examples and describe them in natural language. Early BP benchmarks featured synthetic black-and-white drawings, which might not fully capture the complexity of real-world scenes. Subsequent BP datasets employed real-world images, albeit the represented concepts are identifiable from high-level image features, reducing the task complexity. Differently, the recently released Bongard-RWR dataset aimed at representing abstract concepts formulated in the original BPs using fine-grained real-world images. Its manual construction, however, limited the dataset size to just $60$ instances, constraining evaluation robustness. In this work, we introduce Bongard-RWR+, a BP dataset composed of 5 4005\,400 instances that represent original BP abstract concepts using real-world-like images generated via a vision language model (VLM) pipeline. Building on Bongard-RWR, we employ Pixtral-12B to describe manually curated images and generate new descriptions aligned with the underlying concepts, use Flux.1-dev to synthesize images from these descriptions, and manually verify that the generated images faithfully reflect the intended concepts. We evaluate state-of-the-art VLMs across diverse BP formulations, including binary and multiclass classification, as well as textual answer generation. Our findings reveal that while VLMs can recognize coarse-grained visual concepts, they consistently struggle with discerning fine-grained concepts, highlighting limitations in their reasoning capabilities.
1 Introduction
Abstract visual reasoning (AVR) domain stabinger2021evaluating ; van2021much refers to visual tasks, solving which requires identifying and reasoning about abstract patterns expressed through image-based analogies. Classical AVR tasks and associated benchmark problems include Raven’s Progressive Matrices (RPMs) barrett2018measuring ; zhang2019raven , visual analogy problems hill2019learning ; webb2020learning , and more malkinski2023review . These tasks mainly focus on testing the ability of systematic reasoning and generalisation to new feature distributions. Nevertheless, they typically require supervised training on large-scale datasets. This stands in contrast to human intelligence assessments, which focus on rapid adaptation to novel, potentially never-encountered problems based on prior knowledge.
A distinct alternative to conventional AVR benchmarks is offered by Bongard Problems (BPs), originally proposed in 1970 bongard1970pattern . Each BP consists of two sides, each containing six images, separated according to an abstract rule. The solver’s task is to infer this underlying concept and articulate it in natural language (see Fig. 13). An alternative BP formulation relies on classifying a novel test image (or a pair of test images) to appropriate side(s). Despite its deceptively simple format, the BP setting poses several unique challenges. First, it naturally constitutes a few-shot learning problem fei2006one ; wang2020generalizing – identifying the separating concept requires generalization from just six examples per side. Second, the interpretation of images is inherently contextual linhares2000glimpse . For instance, in Fig. 1(a), a visual feature such as curvature may appear salient on the left side, but only through comparison with the opposite side the true rule, a difference in arrow directions, becomes evident. Third, solving BPs requires integrating visual perception with text generation, demanding bi-modal reasoning capabilities.
Early BPs were created manually, resulting in only a few hundred instances contributed by a number of individuals foundalis2021index . To scale beyond this limitation, Bongard-LOGO nie2020bongard introduced a procedural generation method based on the action-oriented LOGO language, enabling the creation of a large set of problems. While this approach provided sufficient data for training deep models, the resulting BPs were restricted to synthetic black-and-white drawings. Subsequent efforts have shifted BPs into the real-world domain. Bongard HOI jiang2022bongard utilized natural images of human-object interactions. Bongard-OpenWorld wu2024bongardopenworld employed an open-vocabulary of free-form concepts to capture the complexity and ambiguity of real-world scenes. Most recently, Bongard-RWR malkinski2024bongardrwr ; malkinski2024reasoning linked synthetic and real-world domains by using real images to represent abstract concepts found in synthetic BPs, facilitating direct comparison of model performance across both these domains. Notably, malkinski2024reasoning  observed that Bongard-RWR, which focuses on abstract concepts, poses a greater challenge to contemporary models than Bongard HOI and Bongard-OpenWorld, which involve real-world concepts, despite all three using real-world images (see Fig. 2 for an illustration). However, since Bongard-RWR was constructed manually, its small scale limits the robustness and breadth of possible evaluations, highlighting the need for a more scalable approach to dataset construction.
To overcome the scalability limitations of Bongard-RWR, we introduce Bongard-RWR+, a new benchmark featuring real-world representations of selected synthetic BPs (see an overview in Table 1). Unlike Bongard-RWR, which was constructed manually, our approach leverages recent advances in vision language models (VLMs), including Image-to-Text (I2T) and Text-to-Image (T2I) models, to automate dataset creation. For each BP in the original Bongard-RWR, we (1) use Pixtral-12B mistral2024pixtral to describe each image, (2) generate new image descriptions aligned with the matrix concept, (3) employ Flux.1-dev flux2024 to synthesize images from these descriptions, and (4) manually verify that generated images reflect the intended concept (see Fig. 3). This pipeline enables construction of a dataset comprising 5 4005\,400 BPs with the original abstract concepts preserved.
We conduct a comprehensive evaluation on Bongard-RWR+ to systematically assess the reasoning capabilities of state-of-the-art VLMs across a range of problem setups. Specifically, we formulate: (1) binary classification tasks, where the model assigns either a single test image or a pair of images to the correct side(s) of the matrix; (2) a multiclass classification task, where the model selects the concept that best matches the BP matrix from a set of candidates; and (3) a free-form text generation task, in which the model articulates the underlying concept in natural language. Beyond these primary tasks, we perform several ablations to investigate factors influencing model performance. These include examining the effect of model size, comparing color vs. greyscale inputs, contrasting model performance on BPs constructed with real vs. generated images, and varying the number of images per matrix side. Our findings show that, while current VLMs exhibit some capacity to identify high-level, coarse-grained concepts, they consistently struggle with discerning fine-grained concepts, highlighting gaps in their visual reasoning capabilities.
Contributions.
In summary, this paper makes the following contributions:
  1. 1.
    We develop an automated pipeline for generating real-world-like images that express abstract visual concepts.
  2. 2.
    We introduce Bongard-RWR+, a new BP benchmark comprising 5 4005\,400 matrices constructed using this pipeline.
  3. 3.
    We conduct extensive evaluations on the proposed dataset, demonstrating that current VLMs struggle to identify fine-grained visual concepts, revealing important limitations in their AVR abilities.
(c) Bongard-RWR+ (ours)
Figure 1: Bongard Problems. All matrices present the same abstract concept: Left side: Arrows pointing in different directions. Right side: Arrows pointing in the same direction. (a) A manually-designed synthetic BP bongard1970pattern ; foundalis2021index . (b) A manually-designed real-world representation from Bongard-RWR malkinski2024bongardrwr ; malkinski2024reasoning . (c) An automatically generated real-world representation from Bongard-RWR+.
(a) Bongard-LOGO nie2020bongard
(b) Bongard HOI jiang2022bongard
(c) Bongard-OpenWorld wu2024bongardopenworld
Figure 2: Related BP datasets. (a) A synthetic BP from Bongard-LOGO nie2020bongard ; Left: Shapes are the same. Right: Shapes are different. (b) A real-world BP from Bongard HOI jiang2022bongard ; Left: A person driving a car. Right: Not a person driving a car. (c) A real-world BP from Bongard-OpenWorld wu2024bongardopenworld ; Left: The top of a snow-covered mountain. Right: Not the top of a snow-covered mountain. Unlike Bongard-LOGO, which involves synthetic images unfamiliar to VLMs, or Bongard HOI and Bongard-OpenWorld, which focus on coarse-grained concepts, Bongard-RWR+ is designed around abstract concepts expressed through realistic images that require fine-grained visual reasoning.
Table 1: BP datasets. Bongard-RWR+ offers a broad set of matrices that uniquely combine generated real-world-like images with abstract concepts.
Dataset Images Concepts # Matrices # Concepts
Synthetic BPs bongard1970pattern Synthetic Abstract $394$ $388$
Bongard-LOGO nie2020bongard Synthetic Abstract 12 00012\,000 $627$
Bongard HOI jiang2022bongard Real-world Real-world 53 00053\,000 $242$
Bongard-OpenWorld wu2024bongardopenworld Real-world Real-world 1 0101\,010 1 0101\,010
Bongard-RWR malkinski2024reasoning Real-world Abstract $60$ $55$
Bongard-RWR+ (ours) Generated Abstract 5 4005\,400 $49$
2 Related Work
Overview of AVR challenges.
Analogy-making is a cornerstone of human cognition, closely tied to fluid intelligence—the ability to generalize learned skills to novel situations lake2017building . In the visual domain, such tasks present demanding challenges for representation learning chalmers1992high and highlight the need to integrate perception and cognition hofstadter1995fluid , motivating the development of a wide range of AVR benchmarks. Among these, RPMs raven1936mental ; raven1998raven have become a heavily studied problem type hoshen2017iq ; malkinski2024one ; malkinski2022deep ; mandziuk2019deepiq , typically involving a small set of logic-based rules. Datasets like SVRT fleuret2011comparing and those based on BPs jiang2022bongard ; malkinski2024reasoning ; nie2020bongard ; wu2024bongardopenworld introduce more abstract and diverse rules expressed through visual analogies in a few-shot learning framework. Another line of work has proposed generative challenges, such as ARC chollet2019measure and PQA qi2021pqa . Several benchmarks bitton2023vasr ; ichien2021visual ; teney2020v construct AVR tasks using real-world images, enabling evaluation of models pre-trained on large visual datasets. Our proposed dataset builds on this landscape by combining few-shot learning with real-world-like images and supporting both classification and free-form text generation tasks, offering a diverse and challenging testbed for evaluating AVR models.
BP solution methods.
In response to the wide range of AVR challenges, various problem-solving approaches have been developed hernandez2016computer ; malkinski2020multi ; mitchell2021abstraction . Specifically, for BPs, explored methods include program synthesis and inductive logic programming saito1996concept ; sonwane2021using , cognitive architectures foundalis2006phaeaco , Bayesian inference depeweg2018solving ; depeweg2024solving , convolutional networks kharagorgiev2018solving , and meta-learning approaches such as prototypical networks snell2017prototypical , SNAIL mishra2018a , or Meta-Baseline chen2021meta . More recently, LLM-based approaches have emerged, either combining image captioning models with LLMs to process text descriptions wu2024bongardopenworld or leveraging VLMs that inherently handle both modalities malkinski2024reasoning . To establish strong baselines, we conduct experiments using state-of-the-art VLMs. Additionally, prior work has framed BPs in various settings, such as binary classification where test images are assigned to matrix sides nie2020bongard ; kharagorgiev2018solving , or free-form text generation where the model articulates the underlying concept malkinski2024reasoning ; wu2024bongardopenworld . To systematically evaluate model capabilities, we cover a broad range of setups, including image (or description) to side classification, a multiclass concept selection task, and free-form text generation.
Data generation with T2I models.
Generative T2I models have gained significant attention for their broad application. Prior work has applied T2I models to generate synthetic data for supervised learning fan2024scaling ; ravuri2019classification ; sariyildiz2023fake , contrastive representation learning jahanian2022generative ; tian2024learning ; tian2023stablerep , and data augmentation trabucco2024effective . In contrast, we employ a pre-trained T2I model to automatically generate real-world-like images depicting predefined abstract concepts, enabling scalable construction of Bongard-RWR+ instances.
3 Methods
To advance research on abstract reasoning, we introduce Bongard-RWR+, a dataset comprising 5 4005\,400 BPs in its main setting, with additional instances in ablation variants. The dataset supports a range of task formulations, including binary and multiclass classification, and free-form text generation. Each instance is defined as ${\cal BP}=(L,R,T,C)$, where $L=\{L_{1},\ldots,L_{P}\}$ and $R=\{R_{1},\ldots,R_{P}\}$ denote the left and right panel sets (each with $P=6$ images), $T=\{T_{L},T_{R}\}$ contains two test images (one per side), and $C=(C_{L},C_{R})$ specifies the underlying concept in natural language, with $C_{L}/C_{R}$ describing the concept shared by all images on the left / right side, resp.
[…]
Figure 3: Generative pipeline. Starting from a Bongard problem $\mathcal{BP}$ with concept $(C_{L},C_{R})$, the pipeline: (1) describes each image using an I2T model to produce paired positive and negative textual captions $\mathcal{L}^{+}_{i}$ and $\mathcal{L}^{-}_{i}$; (2) augments each positive caption with a T2T model into NN diverse descriptions $\{\mathcal{L}^{+}_{i,j}\}_{j=1}^{N}$ that preserve the underlying concept; (3) generates candidate images for each new description using a T2I model; and (4) involves a human judge to review and filter the generated images. For readability, the figure illustrates the processing flow for the first image from the left matrix side.
3.1 Bongard-RWR+
To generate Bongard-RWR+, we considered each $\cal BP$} from Bongard-RWR. For each image $L_{i}\in L$ and $R_{i}\in R$, we generated paired positive and negative textual descriptions based on the side’s concept: $\text{Describe}(L_{i},C_{L})=(\mathcal{L}^{+}_{i},\mathcal{L}^{-}_{i})$ and $\text{Describe}(R_{i},C_{R})=(\mathcal{R}^{+}_{i},\mathcal{R}^{-}_{i})$, using Pixtral-12B mistral2024pixtral . This I2T model was prompted to produce positive descriptions that faithfully captured the image’s content and negative descriptions designed to steer the T2I model away from depicting the opposite concept. Each positive description was augmented with a Text-to-Text (T2T) model into $N=15$ alternative descriptions reflecting the side’s concept: $\text{Augment}(\mathcal{L}^{+}_{i},C_{L})=\{\mathcal{L}^{+}_{i,j}\}_{j=1}^{N}$ and $\text{Augment}(\mathcal{R}^{+}_{i},C_{R})=\{\mathcal{R}^{+}_{i,j}\}_{j=1}^{N}$. Each generated alternative description, paired with its corresponding negative prompt, was passed to the Flux.1-dev model flux2024 to render $512\times 512$ candidate images: $\text{Render}(\mathcal{L}^{+}_{i,j},\mathcal{L}^{-}_{i})=L_{i,j}$ and $\text{Render}(\mathcal{R}^{+}_{i,j},\mathcal{R}^{-}_{i})=R_{i,j}$. Appendix F describes the employed model prompts. All generated images were manually reviewed to ensure they accurately reflected the intended concept ($C_{L}$ or $C_{R}$) without introducing elements from the opposite side’s concept. Images failing this criterion were discarded.
From the candidate images $\{L^{+}_{i,j}\}$ and $\{R^{+}_{i,j}\}$, we composed $M=10$ left and right sides, denoted $\{L^{\prime}\}_{m=1}^{M}$ and $\{R^{\prime}\}_{n=1}^{M}$, by iteratively selecting subsets that maximized intra-set visual diversity. Each subset $L^{\prime}_{m}\subset\{L^{+}_{i,j}\}$ and $R^{\prime}_{n}\subset\{R^{+}_{i,j}\}$ contained $7$ images ($6$ context images and $1$ test image), chosen iteratively to minimize total pairwise cosine similarity of ViT-L/14 embeddings dosovitskiy2021an . To cover more images and increase dataset diversity, in each iteration we selected a random image from each subset and removed it from the overall image pool in the next iterations. Finally, each left side $L^{\prime}_{m}$ was paired with each right side $R^{\prime}_{n}$ to construct $M^{2}=100$ new BP instances aligned with the original concept. Algorithm details are explained in Appendix A. Using this pipeline, we constructed 5 4005\,400 BPs corresponding to $54$ Bongard-RWR matrices ($100$ new problems per source BP). The $6$ remaining Bongard-RWR matrices were discarded due to difficulties in generating a sufficient number of images faithfully depicting the underlying concepts.
We additionally introduce several dataset variants to support ablation studies. The first variant, Bongard-RWR+/GS, consists of grayscale matrices. Since Bongard-RWR+ concepts are primarily structural, as they originate from black-and-white synthetic BPs, this variant isolates the role of color, which is expected to be non-essential for concept detection. The second variant, Bongard-RWR+/LP ($P=2,\ldots,6$), e.g., denoted Bongard-RWR+/L2 for $P=2$, modifies the subset construction method by omitting the image removal step. In effect, it contains fewer unique images (1 5031\,503 in Bongard-RWR+/L6 vs. 4 1574\,157 in Bongard-RWR+), but exhibits lower average intra-side embedding similarity. Lower similarity implies greater visual diversity within a matrix, meaning the same concept is illustrated through more varied content—for example, the concept "Vertical" may be instantiated as a tree, a building, or a standing figure—potentially making the concept easier to identify. In contrast, high similarity may reflect repeated visual aspects (e.g., several images of trees), which may hint at unrelated concepts like "Nature" or "Green", making the intended concept harder to capture. We consider Bongard-RWR+/LP variants with varying numbers of images per side ($P$), enabling analysis of how the number of demonstrations influences model performance. All variants are derived from the same image pool as the main dataset. Additional details are provided in Appendix A.
[…]
Figure 4: BP formulations. We define six tasks of variable complexity: (a; I1S) assign a single test image to the left or right side; (b; I2S) assign a pair of test images to the respective sides; (c; D1S / d; D2S) use descriptions from an I2T model and classify with a T2T model; (e; CS) select the correct concept index $\widehat{k}$ such that $C_{\widehat{k}}=C^{*}$; (f; CG) generate a natural language description of concept $\widehat{C}$.
3.2 Problem formulations
We cast BP into several task formulations of increasing complexity, as illustrated in Fig. 4. We begin with binary classification settings. In the Image-to-Side (I1S) task, the model classifies a single test image ($T_{L}$ or $T_{R}$) as belonging to either the left or right side. The Images-to-Sides (I2S) variant extends this by requiring the model to assign a pair of test images ($T_{L}$ and $T_{R}$), each from a different class, to their respective sides. To assess the effect of an intermediate image captioning step, we introduce Description-to-Side (D1S) and Descriptions-to-Sides (D2S). These tasks first convert images into natural language descriptions using an I2T model: $\text{Describe}(L_{i})=\mathcal{L}_{i}$, $\text{Describe}(R_{i})=\mathcal{R}_{i}$, $\text{Describe}(T_{s})=\mathcal{T}_{s}$, $\forall i\in\{1,\ldots,P\}$, $\forall s\in\{L,R\}$. The resulting descriptions are passed to a T2T model for a binary classification, either for a single test description (D1S) or a pair (D2S). In practice, I2T and T2T models may be implemented by a single VLM that processes both visual and textual inputs, though T2T can also be realized by text-only models (e.g., LLMs).
Beyond binary classification settings, we introduce the Concept Selection (CS) task, a multiclass classification setup where the model selects the correct concept $C^{*}$ describing the matrix from a candidate set of concepts $\{C_{k}\}_{k=1}^{K}$ ($C^{*}\in\{C_{k}\}$), with $K$ controlling task difficulty. To construct distractors ($\{C_{k}\ |\ C_{k}\neq C^{*}\}$), we sample concepts from other matrices, ensuring that each candidate represents a distinct concept to avoid ambiguity. Finally, we consider the most challenging formulation, Concept Generation (CG), where the model is required to generate a free-form textual description of the concept underlying the BP matrix.
4 Experiments
We evaluate $4$ state-of-the-art open-access VLMs on Bongard-RWR+: InternVL2.5 78B (IVL2.5) chen2024expanding ; chen2024internvl , Qwen2-VL 72B-Instruct (Q2VL) bai2023qwen ; wang2024qwen2 , LLaVA-Next 110B (LLaVA) jiang2023mistral ; liu2024improved ; liu2024visual , and MiniCPM-o 2.6 8B (MCPM) yao2024minicpm ; yu2024rlaif . Main experiments are conducted on the largest available model versions. Smaller variants are covered in ablation studies. All models are evaluated at a fixed decoding temperature of $0.5$, using structured JSON output format enforced via the Outlines decoding backend willard2023efficient . Inference was performed on an internal computing cluster equipped with NVIDIA DGX A100 and H100 nodes. The largest models required up to $24$ hours of inference per task using $4$ H100 GPUs.
We developed a Similarity Classifier (SC) to serve as a non-parametric baseline. For image-based setups (I1S and I2S), it computes ViT-L/14 embeddings dosovitskiy2021an for all matrix images and the test image. For text-based tasks (D1S and D2S), embeddings of image descriptions are obtained with fine-tuned MiniLM minilm2025hugging ; wang2020minilm . Next, for each matrix side, the embedding that is most distant from the test embedding (based on Euclidean distance) is identified. The test input is then assigned to the side with the lower maximum distance (i.e., higher similarity). Additional details are provided in Appendix B.
Figure 5: Concept Selection. Accuracy in the CS task on Bongard-RWR+ for $K\in\{2,4,8,16\}$.
Concept Selection.
We begin with the CS task, where the model selects the correct concept from a candidate set of size $K\in\{2,4,8,16\}$. Increasing $K$ reduces accuracy, confirming the growing difficulty of the task (Fig. 5). InternVL2.5 performs best across all variants, correctly identifying the concept in $91\%$ of cases for $K=2$. However, its accuracy drops to $57\%$ for $K=16$, indicating that even the strongest model struggles when faced with multiple distractors. Meanwhile, both MiniCPM-o 2.6 and LLaVA-Next achieve only $19\%$ for $K=16$, underscoring their limited capacity even in this discriminative setting, and highlighting the utility of the CS task as a benchmark for AVR. Notably, MiniCPM-o 2.6 matches LLaVA-Next’s performance despite a much smaller parameter count (8B vs. 110B), suggesting that size alone is insufficient to excel in the introduced challenge.
Figure 6: Concept Selection. Accuracy in the CS task on Bongard-RWR+ for $K\in\{2,4,8,16\}$.
To better characterize model behavior, we categorized Bongard-RWR+ matrix concepts into $9$ semantic groups reflecting the key factor behind each concept: Size, Position, Count, Branching, Similarity, Contour, Shape, Rotation, and Angle. This grouping follows recent efforts advocating for semantically grounded evaluations to isolate model strengths and weaknesses mitchell2021abstraction ; odouard2022evaluating . The results are shown in Fig. 6. For $K=2$, all models exhibit broadly similar profiles across concept groups. However, for more challenging setups ($K>2$), differences emerge. For instance, InternVL2.5 retains high performance on Shape, Size, and Branching for $K=16$, with accuracy near $75\%$. These categories involve high-level visual attributes that VLMs appear to capture reliably. In contrast, InternVL2.5 accuracy drops below $50\%$ for Contour, Rotation, and Angle. These concepts often rely on subtle visual cues (Contour) or precise spatial relationships (Rotation and Angle), which current models struggle to capture reliably. Overall, performance patterns remain consistent across values of $K$, supporting the utility of such analysis as a diagnostic tool for characterizing model capabilities.
Table 2: *-to-Side(s) classification. Results with I1S, I2S, D1S, and D2S on Bongard-RWR+. Captions in D1S and D2S were produced with InternVL2.5 78B.
I1S I2S D1S D2S
IVL2.5 $0.50$ $0.39$ $0.57$ $0.49$
Q2VL $0.49$ $0.44$ 0.58 $0.42$
LLaVA $0.50$ $0.50$ $0.54$ $0.43$
MCPM $0.48$ $0.45$ $0.51$ $0.41$
DS-R1 N/A N/A $0.57$ 0.56
SC 0.52 0.54 $0.49$ $0.50$
Image(s)-to-Side(s).
Next, we consider the I1S and I2S tasks, where models must assign one or two test images, resp., to the correct matrix side. Notably, in I2S (and D2S), the model performs two independent binary classifications; while the prompt specifies that the test inputs belong to different classes (matrix sides), this constraint is not enforced in the output format, causing some models to incorrectly assign both inputs to the same side. The results are shown in Table 2. For both tasks, VLMs performed at or near chance, with several models falling below random accuracy in I2S. Strikingly, the simple SC baseline outperformed all VLMs, suggesting that the tested models fail to robustly identify the underlying concept that separates semantically the matrix sides. Compared to the CS task, where models can rely on high-level heuristics to eliminate implausible options, I1S and I2S require a deeper understanding of the depicted concept to correctly classify new inputs. These results highlight a fundamental limitation in current VLMs’ AVR capabilities.
Figure 7: Description-to-Side. Accuracy in the D1S task on Bongard-RWR+ for a pair-wise combination of I2T description models (row-wise) and T2T prediction models (column-wise).
Description(s)-to-Side(s).
Given the difficulty of the introduced dataset, we hypothesized that decomposing the problem via an intermediate captioning step may improve model performance. To test this, we generated image descriptions using each of the $4$ VLMs and used these captions as input to the same set of models for solving the side prediction task. Since the prediction task is purely text-based, we also evaluated the DeepSeek-R1 70B (DS-R1) reasoning model guo2025deepseek , which does not support visual input ("N/A" for I1S and I2S in Table 2) but offers strong language understanding. As shown in Fig. 7, final accuracy in the D1S task varies depending on both the captioning (rows) and prediction (columns) models. Descriptions produced by InternVL2.5 yield the highest overall accuracy, with Qwen2-VL also producing relatively strong captions. Interestingly, captions from MiniCPM-o 2.6 lead to better results than those from LLaVA-Next, despite MiniCPM-o 2.6 being much smaller, highlighting its promising image-to-text capabilities. Among prediction models (columns), Qwen2-VL achieves the best average performance, followed by InternVL2.5, indicating relatively strong text-based reasoning abilities of both models. DeepSeek-R1, while known for strong performance on standard benchmarks, ranks third, revealing potential limitations in reasoning over image descriptions. Using captions generated by the best-performing model, InternVL2.5, we report D1S and D2S task performance in Table 2. Compared to I1S and I2S, most models perform above chance in these settings, confirming that the intermediate captioning step helps models better ground their predictions.
(a) Varying model size on Bongard-RWR+
(b) Bongard-RWR+/GS
Figure 8: Concept Selection ablations. (a) Accuracy of models of varying sizes for $K=16$. (b) The impact of using grayscale images for $K\in\{2,4,8,16\}$. Differences w.r.t. color images are shown using lighter colors and are annotated in parentheses above the corresponding bars.
Does AVR performance scale with model size?
Prior works have shown that VLM performance tends to scale with model size hoffmann2022training ; kaplan2020scaling . To examine whether this observation holds for AVR, we evaluated smaller variants of several models in the CS task on Bongard-RWR+. Specifically, we additionally tested 8B, 26B, and 34B variants of InternVL2.5, 7B, 32B, and 72B versions of LLaVA-Next and 7B variant of Qwen2-VL. We excluded additional MiniCPM-o 2.6 variants, as the 8B model already showed weak performance and no larger variants are currently available. As shown in Fig. 8(a), accuracy generally increases with model size. Among the smallest configurations (7B/8B), InternVL2.5 performs best, followed by Qwen2-VL, mirroring the relation observed for their largest versions (70+B). Notably, both InternVL2.5 8B and Qwen2-VL 7B outperformed MiniCPM-o 2.6 8B, underscoring the generally stronger AVR capacity of these two model families.
Is image color necessary for concept recognition?
To assess the role of color in abstract reasoning, we evaluated all $4$ VLMs on Bongard-RWR+/GS, using the CS task. As shown in Fig. 8(b), model performance remains comparable, or even improves, on grayscale images. For instance, LLaVA-Next consistently achieves higher accuracy across all $K$ (e.g. +5+5 p.p. for $K=2$). This outcome aligns with the dataset’s design – the underlying concepts are derived from structural properties of black-and-white synthetic BPs, where color is not semantically relevant. In this context, color may act as a visual distractor that adds additional complexity, especially for lower-performing models.
Figure 9: Functional equivalence of generated images and real-world images. Performance trends in the CS task ($K\in\{2,4,8,16\}$) on Bongard-RWR and Bongard-RWR+/LP for $P=2,3$.
Table 3: I1S and D1S on Bongard-RWR+/LP. Results in I1S and D1S tasks with varying numbers of images per side $P$. Image captions in the D1S task were produced with InternVL2.5 78B.
(a) Image-to-Side (I1S)
$P$ IVL2.5 Q2VL LLaVA MCPM SC
22 $0.51$ $0.51$ $0.49$ $0.50$ 0.54
33 $0.54$ $0.54$ $0.51$ $0.50$ 0.59
44 $0.54$ $0.54$ $0.51$ $0.51$ 0.59
55 $0.54$ $0.55$ $0.51$ $0.52$ 0.58
66 $0.55$ $0.57$ $0.51$ $0.51$ 0.59
(b) Description-to-Side (D1S)
$P$ IVL2.5 Q2VL LLaVA MCPM DS-R1 SC
22 0.56 $0.55$ $0.53$ $0.50$ 0.56 $0.47$
33 $0.60$ $0.58$ $0.56$ $0.51$ 0.62 $0.51$
44 $0.61$ $0.63$ $0.58$ $0.54$ 0.64 $0.51$
55 $0.62$ $0.60$ $0.57$ $0.52$ 0.64 $0.52$
66 $0.67$ $0.67$ $0.61$ $0.53$ 0.70 $0.54$
Are generated images as effective as real ones?
Since Bongard-RWR+ relies on generated images, a key question is whether such images are as effective as real ones for evaluating visual reasoning. To investigate this, we compared model performance on the introduced dataset with that on Bongard-RWR, which consists exclusively of real-world images. However, given that Bongard-RWR includes only $60$ instances, we constructed matrices based on smaller subsets of images per side to expand the dataset. For each BP in Bongard-RWR, we considered all pairs of $P$-element subsets from both matrix sides (analogously as described in Section 3.1), yielding ${6\choose P}^{2}$ new matrices per BP, i.e., $36$, $225$, $400$, $225$, $36$ matrices for $P=1,\ldots,5$, resp. We applied this method to $54$ Bongard-RWR BPs that were used to generate Bongard-RWR+, and then selected $P=2,3$ to ensure sufficient dataset size, and capped the number of sampled matrices per source BP at $100$, resulting in 5 4005\,400 matrices per $P$. We evaluated all $4$ VLMs in the CS task on these Bongard-RWR variants and compared the results to those on Bongard-RWR+/LP, which uses analogous matrix construction approach but with generated images. As shown in Fig. 9, model performance follows similar trends on both datasets – accuracy decreases consistently as $K$ increases, regardless of whether the images are real or generated. This supports the validity of our generation-based approach for capturing the challenges of AVR.
Do models learn from demonstrations?
As $P$ increases, models are presented with more demonstrations of the underlying concept. While this could enhance concept identification, it also increases the amount of visual information to process. Interestingly, Fig. 9 shows that InternVL2.5 and Qwen2-VL benefit from larger $P$ in the CS task, e.g., InternVL2.5’s accuracy improves by $2,3,6,8$ p.p. for $K=2,\ldots,5$, resp. when comparing the results on Bongard-RWR+/L2 with Bongard-RWR+/L3. In contrast, the performance of LLaVA-Next and MiniCPM-o 2.6 shows no consistent improvement, suggesting these models may struggle to leverage additional examples.
To explore this further, we evaluated models on Bongard-RWR+/LP with $P=2,\ldots,6$ using the I1S and D1S tasks. In I1S (Table 3(a)), both InternVL2.5 and Qwen2-VL show a mild upward trend with increasing $P$, with Qwen2-VL achieving the best accuracy of $57\%$ at $P=6$, and InternVL2.5 close behind at $55\%$. In D1S (Table 3(b)), accuracy generally improves with $P$, showing that the models can effectively utilize additional demonstrations for concept identification in a text-based setting. Notably, DeepSeek-R1 outperforms others for every $P>2$ (and performs on-par with InternVL2.5 for $P=2$), demonstrating its strength in this setting. Additionally, D1S scores are consistently higher than I1S scores, showing that the employed models benefit from the explicit image captioning step. Nevertheless, performance remains modest across both tasks, in I1S being clearly inferior to the SC baseline, which highlights a persistent gap in AVR among the contemporary models.
Finally, model accuracy on Bongard-RWR+/LP exceeds that on Bongard-RWR+ (e.g., InternVL2.5 scores $50\%$ and $57\%$ in I1S and D1S, resp. on Bongard-RWR+ (cf. Table 2), but reaches $55\%$ and $67\%$ on Bongard-RWR+/L6). This gap stems from the differences in dataset construction, which we analyze in detail in Appendix B.
Concept Generation.
We also assessed the ability of the models to generate free-form descriptions of underlying concepts in the CG task. Models were prompted to describe the concept behind each matrix, using either raw image inputs or InternVL2.5’s image captions. The generated responses were evaluated using standard NLP metrics: BLEU papineni2002bleu , METEOR banerjee2005meteor , $\text{ROUGE}_{L}$ lin2004rouge , CIDEr lin2004rouge , and BERTScore zhang2020bertscore . The results, reported in Appendix B, show that the models consistently achieved low scores on all metrics, with no clear best-performing model. Manual inspection of selected outputs further confirmed that current models generally struggle to articulate the BP concepts, highlighting this task as a particularly challenging direction for future research.
5 Conclusions
We introduced Bongard-RWR+, a large-scale dataset for evaluating AVR capabilities using generated real-world-like images that depict abstract concepts originating from classical BPs. The dataset supports a diverse set of BP task formulations, including multiclass concept selection, binary side classification, and free-form text generation. Our results show that while current VLMs demonstrate certain ability to identify coarse-grained concepts, they consistently struggle with fine-grained concept recognition. Although explicit image captioning and increased visual demonstrations can improve performance, even the strongest models fall short in more demanding setups, in particular, the unconstrained textual answer generation.
Limitations.
Our dataset generation pipeline still requires human oversight to ensure adherence to intended concepts; advancing generative modeling is needed to improve scalability. Our evaluation includes selected VLMs; broader community involvement is essential to benchmark additional methods on Bongard-RWR+. A systematic review of AVR performance across datasets, such as VCog-Bench cao2024visual , MindSet biscione2024mindset , or WAIS-IV galatzer2024cognitive , could further contextualize model capabilities. We elaborate on these promising research directions in Appendix C.
Appendix A Dataset variants
[…]
2:Output: Generated images $G=(GL$, $GR)$
[…]
4:
5:for $(L_{k},R_{k},T_{k},C_{k})\in{\cal RWR}$ do
6:  $L_{k}\leftarrow L_{k}\cup T_{k,L}$
7:  $R_{k}\leftarrow R_{k}\cup T_{k,R}$
8:  for $S_{k}\in\{L_{k},R_{k}\}$ do
9:   for $S_{k,i}\in S_{k}$ do
10:    $({\cal S}_{k,i}^{+},{\cal S}_{k,i}^{-})\leftarrow\text{Describe}(S_{k,i},C_{S_{k}})$
11:    $\{{\cal S}_{k,i,j}^{+}\}_{j=1}^{N}\leftarrow\text{Augment}({\cal S}_{k,i}^{+},C_{S_{k}})$
12:
13:    for j=1,,Nj=1,...,N do
14:     $GS_{k,i,j}\leftarrow\text{Render}({\cal S}_{k,i,j}^{+},{\cal S}_{k,i}^{-})$
15:    end for
16:   end for
17:  end for
18:end for
Algorithm 1 The generation of Bongard-RWR+ images.
Image generation.
Algorithm 1 outlines the process of generating Bongard-RWR+ images. Each BPk $(L_{k},R_{k},T_{k},C_{k})$ from Bongard-RWR is processed by iterating over both matrix sides, $L_{k}$ and $R_{k}$. Test images $T_{k}$ are processed based on their corresponding side. For each image $S_{k,i}$ on a given side $S_{k}$ (either $L_{k}$ or $R_{k}$), a positive ($\mathcal{S}_{k,i}^{+}$) and a negative ($\mathcal{S}_{k,i}^{-}$) textual description are generated based on the side’s concept $C_{S_{k}}$ (either $C_{L_{k}}$ or $C_{R_{k}}$). The positive description is then augmented $N=15$ times to produce a diverse set of semantically consistent variations $\mathcal{S}_{k,i,j}^{+},j=1,\ldots,N$. Each augmented description is paired with the corresponding negative description and passed to a T2I model to generate a new image $GS_{k,i,j}$ (either $GL_{k,i,j}$ or $GR_{k,i,j}$). All generated images were reviewed manually by a human expert, and those that did not faithfully reflect the intended concept were discarded.
1:Input: Generated images $G=(GL,GR)$, Boolean image removal flag $F$, Number of images per side $P$
[…]
5:
6:for $GL_{k},GR_{k}\in G$ do
7:  for $GS_{k}\in\{GL_{k},GR_{k}\}$ do
8:   $E_{k}\leftarrow\text{GetImageEmbedings}(GS_{k})$
9:   $GS_{k}^{*}\leftarrow\emptyset$
10:   $\{\mathbf{GS_{k,l}},\mathbf{E_{k,l}}\}\leftarrow\text{GetAllSubsets}(GS_{k},E_{k},P)$
11:   for $m=1,\dots,M$ do
12:    // Find the most diverse subset
13:    $\ell\leftarrow\operatorname*{arg\,min}_{l=1,\dots,|\mathbf{E_{k}}|}\text{max}\{\mathbf{E_{k,l,i}}\cdot\mathbf{E_{k,l,j}}\ |\ i,j=1,\dots,P\ \land\ i\neq j\}$
14:    $GS_{k}^{*}\leftarrow GS_{k}^{*}\cup\mathbf{GS_{k,\ell}}$
15:
16:    if $F$ then
17:     // Find the most redundant image and remove it from the overall image pool
18:     $p\leftarrow\operatorname*{arg\,max}_{i=1,\dots,P}\{\mathbf{E_{k,\ell,i}}\cdot\mathbf{E_{k,\ell,j}}\ |\ j=1,\dots,P\ \land\ i\neq j\}$
19:     $GS_{k}\leftarrow GS_{k}-\{\mathbf{GS_{k,\ell,p}}\}$
20:     $E_{k}\leftarrow E_{k}-\{\mathbf{E_{k,\ell,p}}\}$
21:     $\{\mathbf{GS_{k}},\mathbf{E_{k}}\}\leftarrow\text{RemoveSubsetsContainingImage}(\mathbf{GS_{k}},\mathbf{E_{k},\mathbf{GS_{k,\ell,p}}})$
22:     // Exit early if there is not enough images to construct the next subset
23:     if $|GS_{k}|<P$ then
24:      break
25:     end if
26:     // Exit early if the images become too similar
27:     $s\leftarrow\text{Mean}(\{\frac{E_{k,i}\cdot E_{k,j}}{\|E_{k,i}\|\|E_{k,j}\|}\ |\ i,j=1,\dots,|E_{k}|\ \land\ i\neq j\})$
28:     if $s\geq 0.85$ then
29:      break
30:     end if
31:    end if
32:   end for
33:  end for
34:
35:  ${\cal RWR^{+}}\leftarrow{\cal RWR^{+}}\cup(GL_{k}^{*}\times GR_{k}^{*})$
36:end for
Algorithm 2 Bongard-RWR+ matrix construction.
Bongard-RWR+ construction.
[…]
Bongard-RWR+/GS construction.
We constructed a grayscale variant of Bongard-RWR+, denoted Bongard-RWR+/GS, by converting all images in the original set to grayscale. The overall matrix construction procedure remained unchanged. Representative examples are shown in Fig. 14.
Bongard-RWR+/LP construction.
We also constructed Bongard-RWR+/LP, a dataset variant in which the image removal step is disabled (i.e., the flag $F$ in Algorithm 2 is set to false). Without this removal step, the greedy subset selection procedure for building $GS_{k}^{*}$ often draws from a smaller set of unique images within $GS_{k}$. In effect, matrices in Bongard-RWR+/LP exhibit lower average intra-side embedding similarity but are composed of fewer distinct images. The increased visual diversity within each matrix can facilitate recognition by presenting concepts in a broader range of views. However, because the same images are reused more frequently across matrices, there is overall less image-level diversity compared to Bongard-RWR+. We constructed dataset variants with $P=2,\ldots,6$, denoted Bongard-RWR+/L2, Bongard-RWR+/L3, etc. Example BPs from Bongard-RWR+/LP are shown in Figs. 15 – 19.
Construction of Bongard-RWR+/TVT and Bongard-RWR+/TVT-Large.
We also developed Bongard-RWR+/TVT, a dataset variant in which BPs are divided into train, validation, and test splits (TVT), enabling the evaluation of supervised approaches. To construct these splits, we first consider the six images on each side of a Bongard-RWR matrix, and the corresponding seventh test image. Images 1 – 4, along with the side’s test image, are grouped into a shared pool used to construct BP contexts for all three splits and to define test targets for the training split. Images 5 and 6 are exclusively allocated as test images for the validation and test splits, resp. This partitioning strategy ensures relatively high image diversity in the training set while enabling evaluation on out-of-distribution images in the validation and test splits. Bongard-RWR+/TVT comprises 12 15012\,150 matrices (7 2907\,290 / 1 2151\,215 / 3 6453\,645 in train / validation / test splits); we additionally provide an extended variant, Bongard-RWR+/TVT-Large, consisting of 86 40086\,400 BPs (51 84051\,840 / 8 6408\,640 / 25 92025\,920 in train / validation / test splits).
Figure 10: Solving Bongard-RWR+/GS with I1S and I2S. The impact of using grayscale images on accuracy in the I1S (left) and I2S (right) tasks.
Figure 11: Functional equivalence of generated images and real-world images. Performance trends in the I1S task on Bongard-RWR and Bongard-RWR+/LP for $P=2,3,4$.
Appendix B Extended results
B.1 Is image color necessary for concept recognition?
To extend the grayscale image analysis from the CS task, we evaluated model performance on Bongard-RWR+/GS with the I1S and I2S tasks. As shown in Fig. 10, grayscale images led to improved accuracy in certain cases, likely by encouraging the model to focus on structural cues. Nevertheless, performance remained close to the random guess level across both tasks, reiterating their inherent difficulty.
B.2 Are generated images as effective as real ones?
We evaluated model performance on Bongard-RWR and Bongard-RWR+/LP in the I1S task for $P=2,3,4$. As shown in Fig. 11, results are comparable across both datasets, regardless of whether the model processes real-world or generated images. This finding is consistent with the experiment on the CS task presented in the main paper and reinforces our claim that generated images are equally effective as real-world images for probing visual reasoning capabilities of VLMs.
Table 4: Concept Selection on Bongard-RWR+/LP for $K=16$.
$P$ IVL2.5 Q2-VL LLaVA MCPM
22 $0.46$ $0.37$ $0.18$ $0.18$
33 $0.54$ $0.41$ $0.19$ $0.17$
44 $0.46$ $0.39$ $0.19$ $0.18$
55 $0.58$ $0.39$ $0.19$ $0.19$
66 $0.62$ $0.41$ $0.20$ $0.18$
B.3 Do models learn from demonstrations?
We analyzed how the number of images per side ($P$), representing concept demonstrations, affects model performance. Extending the I1S and D1S analysis in the main paper, we evaluated model accuracy on Bongard-RWR+/LP using the CS task across varying values of $P$. As shown in Table 4, performance generally improves with more demonstrations, particularly for stronger models such as InternVL2.5, which achieves its best result at $P=6$. However, not all models exhibit this trend – Qwen2-VL performs similarly for $P=3$ and $P=6$, while MiniCPM-o 2.6 achieves its peak performance for $P=5$. These results suggest that while additional demonstrations can be beneficial, some models may not be able to fully exploit them.
Table 5: Concept Generation: Image-based on Bongard-RWR+.
$\text{BLEU}_{1}$ $\text{BLEU}_{2}$ METEOR $\text{ROUGE}_{L}$ CIDEr $P_{\text{BERT}}$ $R_{\text{BERT}}$ $F_{\text{BERT}}$
InternVL2.5 78B $0.063$ 0.011 $0.056$ $0.159$ 0.017 $0.114$ 0.049-0.049 $0.027$
Qwen2-VL 72B $0.048$ $0.004$ $0.043$ $0.121$ $\underline{0.007}$ $0.066$ 0.102-0.102 0.025-0.025
LLaVA-Next 110B $\underline{0.071}$ $0.008$ $\underline{0.061}$ $\underline{0.174}$ $0.006$ 0.128 $\underline{-0.037}$ $\underline{0.040}$
MiniCPM-o 2.6 8B 0.077 $\underline{0.008}$ 0.063 0.181 $0.004$ $\underline{0.127}$ -0.035 0.041
Table 6: Concept Generation: Image-based on Bongard-RWR+/L6.
$\text{BLEU}_{1}$ $\text{BLEU}_{2}$ METEOR $\text{ROUGE}_{L}$ CIDEr $P_{\text{BERT}}$ $R_{\text{BERT}}$ $F_{\text{BERT}}$
InternVL2.5 78B $0.060$ 0.011 $0.053$ $0.149$ 0.021 $0.111$ 0.045-0.045 $0.028$
Qwen2-VL 72B $0.042$ $0.005$ $0.039$ $0.110$ $\underline{0.008}$ $0.046$ 0.118-0.118 0.042-0.042
LLaVA-Next 110B $\underline{0.072}$ $0.007$ $\underline{0.061}$ $\underline{0.172}$ $0.008$ $\underline{0.133}$ $\underline{-0.031}$ $\underline{0.045}$
MiniCPM-o 2.6 8B 0.076 $\underline{0.009}$ 0.064 0.182 $0.005$ 0.144 -0.024 0.054
Table 7: Concept Generation: Text-based on Bongard-RWR+.
$\text{BLEU}_{1}$ $\text{BLEU}_{2}$ METEOR $\text{ROUGE}_{L}$ CIDEr $P_{\text{BERT}}$ $R_{\text{BERT}}$ $F_{\text{BERT}}$
InternVL2.5 78B $0.122$ $\underline{0.021}$ $0.063$ $0.158$ $0.017$ $0.142$ $\underline{0.073}$ $\underline{0.107}$
Qwen2-VL 72B 0.132 0.021 0.066 $0.161$ $\underline{0.018}$ $\underline{0.145}$ 0.090 0.117
LLaVA-Next 110B $0.111$ $0.018$ $0.064$ $\underline{0.165}$ $0.013$ 0.147 $0.062$ $0.103$
MiniCPM-o 2.6 8B $0.105$ $0.019$ $\underline{0.064}$ 0.170 0.019 $0.141$ $0.033$ $0.085$
DeepSeek-R1 70B $\underline{0.129}$ $0.013$ $0.060$ $0.138$ $0.015$ $0.069$ $0.050$ $0.060$
Table 8: Concept Generation: Text-based on Bongard-RWR+/L6.
$\text{BLEU}_{1}$ $\text{BLEU}_{2}$ METEOR $\text{ROUGE}_{L}$ CIDEr $P_{\text{BERT}}$ $R_{\text{BERT}}$ $F_{\text{BERT}}$
InternVL2.5 78B $0.123$ $\underline{0.023}$ $0.065$ $0.162$ $\underline{0.019}$ 0.154 $\underline{0.087}$ $\underline{0.120}$
Qwen2-VL 72B 0.132 $0.023$ 0.066 $0.160$ $0.018$ $\underline{0.153}$ 0.095 0.124
LLaVA-Next 110B $0.114$ $0.020$ $0.064$ $\underline{0.164}$ $0.012$ $0.149$ $0.070$ $0.108$
MiniCPM-o 2.6 8B $0.110$ 0.024 $\underline{0.066}$ 0.169 0.022 $0.146$ $0.046$ $0.094$
DeepSeek-R1 70B $\underline{0.128}$ $0.015$ $0.060$ $0.134$ $0.018$ $0.072$ $0.052$ $0.062$
Table 9: Concept Generation: Image-based on Bongard-RWR+ vs. Bongard-OpenWorld. Comparison of concept descriptions generated from images in Bongard-RWR+ (using MiniCPM-o 2.6 $8\text{B}$) and Bongard-OpenWorld (using GPT-4V (wu2024bongardopenworld, , Appendix E, Table 4)).
$\text{BLEU}_{1}$ $\text{BLEU}_{2}$ METEOR $\text{ROUGE}_{L}$ CIDEr
Bongard-RWR+ $0.077$ $0.008$ $0.063$ $0.181$ $0.004$
Bongard-OpenWorld $0.190$ $0.073$ $0.111$ $0.188$ $0.527$
Table 10: Concept Generation: Text-based on Bongard-RWR+ vs. Bongard-OpenWorld. Comparison of concept descriptions generated from image captions in Bongard-RWR+ (using Qwen2-VL $72\text{B}$ for prediction and InternVL2.5 $78\text{B}$ for captioning) and Bongard-OpenWorld (using ChatGPT for prediction and BLIP-2 w/ Fine-tuning for captioning (wu2024bongardopenworld, , Appendix E, Table 4)).
$\text{BLEU}_{1}$ $\text{BLEU}_{2}$ METEOR $\text{ROUGE}_{L}$ CIDEr
Bongard-RWR+ $0.132$ $0.021$ $0.066$ $0.161$ $0.018$
Bongard-OpenWorld $0.441$ $0.292$ $0.222$ $0.417$ $1.714$
Table 11: Baseline performance. Test accuracy of supervised learning models on Bongard-RWR+/TVT and Bongard-RWR+/TVT-Large, using image and text-based input representations. Each experiment was repeated $10$ times with different seeds; the table reports mean $\pm$ std.
Bongard-RWR+/TVT Bongard-RWR+/TVT-Large
Image Text Image Text
MLP $0.49\pm 0.00$ $0.49\pm 0.00$ $0.47\pm 0.01$ $0.49\pm 0.00$
WReN $0.49\pm 0.01$ $0.48\pm 0.01$ $0.48\pm 0.01$ $0.48\pm 0.01$
SNAIL $0.48\pm 0.01$ $0.51\pm 0.01$ $0.48\pm 0.01$ $0.48\pm 0.01$
B.4 Additional baselines
We evaluated the performance of supervised learning methods on the I1S and D1S tasks. The models operated on embeddings produced by pre-trained models. For I1S, raw images were encoded using ViT-L/14 dosovitskiy2021an , while for D1S, image captions produced by InternVL2.5 were embedded using fine-tuned MiniLM minilm2025hugging ; wang2020minilm . We considered an MLP with a single hidden layer applied to concatenated embeddings, Wild Relation Network (WReN) barrett2018measuring , which processes embedding pairs and aggregates them via summation, and SNAIL mishra2018a , an attention-based meta-learner. As shown in Table 11, all models performed near the random guess level on both Bongard-RWR+/TVT and Bongard-RWR+/TVT-Large, demonstrating that the proposed BPs present a significant challenge not only for VLMs but also for standard supervised learning approaches.
B.5 Performance on Concept Generation task
Tables 5 – 8 evaluate model performance on the CG task, where models are prompted to produce free-form descriptions of the underlying concept for each matrix. Inputs were either raw images or image captions produced by InternVL2.5. Experiments were conducted on both Bongard-RWR+ and Bongard-RWR+/L6 datasets. We assess generated descriptions using standard NLP metrics: BLEU ($\text{BLEU}_{1}$ and $\text{BLEU}_{2}$papineni2002bleu , METEOR banerjee2005meteor , $\text{ROUGE}_{L}$ lin2004rouge , CIDEr lin2004rouge , and BERTScore (precision $P_{\text{BERT}}$, recall $R_{\text{BERT}}$, F1 $F_{\text{BERT}}$zhang2020bertscore . For BERTScore, we apply baseline rescaling [^1] https://github.com/Tiiiger/bert_score/blob/master/journal/rescale_baseline.md and the DeBERTa [^2] https://huggingface.co/microsoft/deberta-xlarge-mnli model he2021deberta , which has strong alignment with human evaluations. Higher scores indicate better performance across all metrics.
MiniCPM-o 2.6 consistently performs best in image-based setups (Tables 5 – 6), ranking first in $5$ of $8$ metrics on Bongard-RWR+ and $6$ of $8$ on Bongard-RWR+/L6, though differences with other models are small. In text-based setups (Tables 7 – 8), Qwen2-VL leads in $5$ of $8$ metrics on Bongard-RWR+ and $4$ of $8$ on Bongard-RWR+/L6, again with relatively small gaps to competing models. We further compare top-performing models on Bongard-RWR+ with the best results reported for Bongard-OpenWorld (wu2024bongardopenworld, , Appendix E, Table 4), as shown in Tables 9 and 10. Regardless of input representation, models achieve significantly higher scores on Bongard-OpenWorld than on Bongard-RWR+, indicating that the concepts in our dataset are more challenging to recognize even for state-of-the-art models.
Table 12: Relative difficulty of Bongard-RWR+/L6 vs. Bongard-RWR+. Each entry reports accuracy on Bongard-RWR+/L6, with the absolute change vs. Bongard-RWR+ shown in parentheses.
Image Text
CS, $K=16$ I1S I2S D1S D2S
InternVL2.5 $78\text{B}$ $\textbf{0.62}\ (+0.05)$ 0.55(+0.05)0.55\ (+0.05) 0.48(+0.09)0.48\ (+0.09) 0.67(+0.10)0.67\ (+0.10) 0.64(+0.15)0.64\ (+0.15)
Qwen2-VL $72\text{B}$ 0.41(+0.03)0.41\ (+0.03) 0.57(+0.08)0.57\ (+0.08) 0.52(+0.08)0.52\ (+0.08) 0.67(+0.09)0.67\ (+0.09) 0.56(+0.14)0.56\ (+0.14)
LLaVA-Next $110\text{B}$ 0.20(+0.01)0.20\ (+0.01) 0.51(+0.01)0.51\ (+0.01) 0.49(0.01)0.49\ (-0.01) 0.61(+0.07)0.61\ (+0.07) 0.52(+0.09)0.52\ (+0.09)
MiniCPM-o 2.6 $8\text{B}$ 0.18(0.01)0.18\ (-0.01) 0.51(+0.03)0.51\ (+0.03) 0.51(+0.06)0.51\ (+0.06) 0.53(+0.02)0.53\ (+0.02) 0.45(+0.04)0.45\ (+0.04)
DeepSeek-R1 $70\text{B}$ N/A N/A N/A $\textbf{0.70}\ (+0.13)$ $\textbf{0.72}\ (+0.16)$
Similarity Classifier N/A $\textbf{0.59}\ (+0.07)$ $\textbf{0.68}\ (+0.14)$ 0.54(+0.05)0.54\ (+0.05) 0.59(+0.09)0.59\ (+0.09)
Figure 12: Model performance on the CS, I1S and D1S tasks across concept groups on Bongard-RWR+. The presented CS results are for $K=4$.
B.6 Assessing difficulty of Bongard-RWR+ variants
We compared model performance on a range of tasks using both Bongard-RWR+ and Bongard-RWR+/L6, with the results shown in Table 12. Overall, the models consistently perform better on Bongard-RWR+/L6 across both image- and text-based strategies. In particular, DeepSeek-R1 presents notable gains in text-based setups, achieving $70\%$ accuracy with D1S and $72\%$ using D2S. We attribute these improvements to the greedy matrix construction strategy used in Bongard-RWR+/L6, which prioritizes subsets that minimize intra-side embedding similarity. This encourages the selection of more visually diverse images to represent a concept, making the underlying shared aspects more distinguishable and thus easier to recognize for the models.
Table 13: Top 55 correctly recognized concepts for each model using the I1S task on Bongard-RWR+.
Model Accuracy Concept Group Left-side Rule Right-side Rule
IVL2.5 $0.64$ Contour Shading thicker on the right side Shading thicker on the left side
$0.61$ Similarity Three identical elements Four identical elements
$0.61$ Shape A circle No circle
$0.59$ Count Three parts Five parts
$0.58$ Angle Acute angle No acute angle
Q2-VL $0.65$ Count Empty picture Not empty picture
$0.62$ Angle A sharp projection No sharp projection
$0.61$ Angle Convex figures Nonconvex figures
$0.59$ Position Axes of symmetry No axes of symmetry
$0.59$ Branching The chain does not branch The chain branches
LLaVA $0.55$ Angle A sharp projection No sharp projection
$0.55$ Rotation An acute angle directed inward No angle directed inward
$0.55$ Similarity All figures of the same color Figures of different colors
$0.54$ Position Points inside the figure outline are on a straight line Points inside the figure outline are not on a straight line
$0.54$ Similarity Figures are similar Figures are not similar
MCPM $0.61$ Similarity All figures of the same color Figures of different colors
$0.60$ Count One figure Two figures
$0.58$ Similarity Figures are similar Figures are not similar
$0.53$ Angle Convex figures Nonconvex figures
$0.53$ Branching There are no side branches of the second order There are side branches of the second order
Table 14: Top 55 correctly recognized concepts for each model using the D1S task on Bongard-RWR+.
Model Accuracy Concept Group Left-side Rule Right-side Rule
IVL2.5 $0.98$ Count One figure Two figures
$0.93$ Count Three parts Four parts
$0.86$ Shape Triangles Circles
$0.84$ Size Large figures Small figures
$0.78$ Count Three parts Five parts
Q2-VL $0.96$ Count Three parts Four parts
$0.95$ Count One figure Two figures
$0.85$ Shape Triangles Circles
$0.83$ Size Large figures Small figures
$0.81$ Shape Polygons Curvilinear figures
LLaVA $0.86$ Count Three parts Four parts
$0.85$ Count One figure Two figures
$0.81$ Shape Polygons Curvilinear figures
$0.78$ Size Large figures Small figures
$0.71$ Count Three parts Five parts
MCPM $0.80$ Count Three parts Four parts
$0.73$ Count One figure Two figures
$0.59$ Count Three parts Five parts
$0.59$ Similarity Three identical elements Four identical elements
$0.57$ Similarity All figures of the same color Figures of different colors
DS-R1 $1.00$ Count One figure Two figures
$0.98$ Count Three parts Four parts
$0.85$ Shape Polygons Curvilinear figures
$0.84$ Shape Triangles Circles
$0.78$ Size Large figures Small figures
B.7 Qualitative analysis
Tables 13 and 14 present the top $5$ correctly recognized concepts per model for the I1S and D1S tasks, resp. In the image-based setup, we observe a diverse distribution of top concepts across the models, with little overlap. This variation suggests that each model relies on distinct reasoning strategies when processing visual input. For example, InternVL2.5 shows some capacity for identifying rules involving contour, similarity, and shape, whereas Qwen2-VL performs better on concepts related to count and angle. In contrast, the text-based setup yields more consistent results across models. Counting-related concepts dominate the top-performing examples, followed by shape and size.
The contrast between I1S and D1S results highlights modality-specific strengths, as the models tend to prefer different concept groups depending on whether they are reasoning over images or text. These trends are further illustrated in Fig. 12, which shows average accuracy per concept group. Most models, InternVL2.5 and Qwen2-VL in particular, exhibit performance spikes in size, count, and shape groups with the D1S task compared to I1S. This discrepancy indicates a weak connection between visual and textual processing pathways in multimodal systems and suggests that improving the integration between these modalities could further boost model performance via knowledge transfer.
Appendix C Limitations
Our dataset generation pipeline requires manual verification to ensure that generated images accurately reflect the intended concepts. During initial experiments, we identified limitations in the ability of T2I models to reliably render certain abstract or fine-grained visual concepts. For instance, the concept "there are (no) inside figures of the second order" from Bongard-RWR could not be accurately represented by the T2I model, preventing us from including it in Bongard-RWR+. We believe future advances in accurate representation of fine-grained concepts in generative models will be critical for improving scalability of datasets like Bongard-RWR+.
Our main experiments evaluated the ability of VLMs to solve BPs, alongside selected supervised learning baselines described in Appendix B. The results highlight clear limitations in the current models’ capacity for abstract visual reasoning, especially when dealing with fine-grained concepts, raising the need for more sophisticated approaches. In particular, multimodal reasoning models that integrate visual and textual information may offer a promising direction, analogous to how reasoning models like DeepSeek-R1 operate in the text domain. We envision the introduced datasets as valuable benchmarks for tracking progress in AVR capabilities.
Multiple recent studies confirm the limitations of VLMs highlighted in our work. MindSet: Vision biscione2024mindset tests models on $30$ psychological findings inspired by human visual perception, revealing fundamental differences between human and machine vision. VCog-Bench cao2024visual evaluates models on abstract and commonsense visual reasoning as well as visual question answering, showing that VLMs consistently struggle with multi-image reasoning tasks. Evaluations on the WAIS-IV test indicate that VLMs underperform in perceptual reasoning tasks galatzer2024cognitive . A systematic review of AVR performance across such benchmarks could provide a more comprehensive understanding of current model capabilities and limitations.
Appendix D Broader impacts
Our findings highlight that state-of-the-art VLMs continue to struggle with tasks that are relatively straightforward for humans. As public discourse increasingly focuses on the capabilities and risks of AI systems, benchmarks such as Bongard-RWR+ play a crucial role in dispelling misconceptions about current model capabilities. Our release of code and datasets under an open license promotes open research and lowers the barrier of entry for researchers to evaluate and improve their methods. However, we also acknowledge potential risks. An effective AVR solver could potentially be misused in online or remote IQ assessments, allowing individuals to inflate their scores. Such misuse may lead to unfair advantages in job recruitment or other competitive settings where cognitive ability plays a role. This underscores the importance of raising public awareness about both the capabilities and potential applications of such tools.
Appendix E Data and Code
All datasets developed in this work are publicly available via the HuggingFace Datasets platform [^3] https://huggingface.co/collections/spawlonka/bongard-rwr-plus-68252881607162f6bc884efa, released under the CC BY 4.0 License. This licensing is compatible with the terms of the Flux.1-dev model [^4] https://huggingface.co/black-forest-labs/FLUX.1-dev/blob/main/LICENSE.md, which was used to generate the images. Code for dataset generation and experimental evaluation is available on GitHub [^5] https://github.com/pavonism/bongard-rwr-plus under the MIT License.
Appendix F Model prompts
Prompts $F$, and F correspond to the Describe, Augment, and Render steps in Algorithm 1, resp. Prompts F, F, F, F, and $F$ are used for the CS, I1S, D1S, I2S, and D2S tasks. Prompt F outlines the shared task introduction for both I1S and D1S, while Prompt F provides the common introduction for I2S and D2S. Prompts include $2$ illustrative examples to help models understand the task format. Importantly, these examples use different class labels to avoid biasing the model toward always selecting the same answer as in the examples. When applicable, models are instructed to respond in a structured JSON format. The validity of these outputs is enforced using the Outlines decoding backend willard2023efficient .
Prompt 1: Generate both positive and negative image descriptions. You are given an image and a general concept. Your task is to refine the concept into two concise, visually descriptive prompts: one that aligns with the image (positive prompt) and one that contrasts with it (negative prompt). Focus on making each prompt specific, clearly grounded in the image, and reflective of the core idea. You don’t need to match every detail—just convey the main visual concept. ### Example: Image: Human legs wearing socks with vertical lines Concept: Vertical lines. Positive prompt: Socks with vertical lines Negative prompt: Socks with horizontal lines Now, it’s your turn: Concept: {concept} ### Instructions: 1. Generate a positive and negative prompt based on the provided image and concept. 2. Answer using the following format: Positive prompt: Negative prompt:
Prompt 2: Augment a positive image description. You are tasked with modifying a given prompt for a diffusion model. Your primary objective is to preserve the specified **concept** while altering unrelated details or environments. You are encouraged to create diverse, creative, and unique augmentations that stay true to the concept but introduce variety in interpretation. ### Example: Prompt: "An empty white bowl with a thin black rim placed on a solid blue background." Concept: "Empty picture" Output Augmentations: [ "An empty red plate on a wooden table.", "A clear glass cup sitting on a marble countertop.", "A white ceramic vase on a patterned fabric." ] Now, it’s your turn: Prompt: {prompt} Concept: {concept} ### Instructions: 1. Generate **{n_augmentations} unique augmentations** for the provided prompt. 2. Output the results in the following JSON string array format: [ "<augmented_prompt_1>", "<augmented_prompt_2>", ... ] Ensure that each augmentation aligns with the concept and introduces creative variations in other details.
Prompt 3: Positive prompt for rendering new images. Negative prompt was provided unchanged. Sharp, high quality image of <positive prompt>.
Prompt 4: Concept Selection (CS) task prompt. You are a vision understanding module (an expert) designed to provide short, clear and accurate answers. The goal in solving a Bongard Problem is to identify a concept that differentiates the left and right matrix sides. You will receive a list of concepts. Only one of them correctly describes the image. Select the concept that correctly describes the image. Respond in JSON using the following format. EXAMPLE START REQUEST { "concepts": [ { "left": "Parallel lines", "right": "Perpendicular lines", "label": 1 }, { "left": "Negative slope", "right": "Positive slope", "label": 2 } ] } RESPONSE { "explanation": "The image presents several lines that don’t intersect.", "label": 1 } EXAMPLE END EXAMPLE START REQUEST { "concepts": [ { "left": "Two shapes", "right": "Six shapes", "label": 1 }, { "left": "A cycle", "right": "No cycle", "label": 2 }, { "left": "Similar shapes", "right": "Different shapes", "label": 3 }, { "left": "Small variance", "right": "Large variance", "label": 4 } ] } RESPONSE { "explanation": "All images on the left side feature circles of similar size, while the images on the right side present triangles, squares, and lines.", "label": 3 } EXAMPLE END REQUEST <request> RESPONSE
Prompt 5: Shared task introduction used in both I1S and D1S settings. You are a vision understanding module (an expert) designed to provide short, clear and accurate answers. The goal in solving a Bongard Problem is to classify a test image to the corresponding class. There are two classes: LEFT and RIGHT. All images belonging to the LEFT class represent a common, shared concept which is not present in any image from the RIGHT class, and vice versa - all images belonging to the RIGHT class represent a common, shared concept which is not present in any image from the LEFT class.
Prompt 6: Image-to-Side (I1S) task prompt. Solve the provided Bongard Problem. Respond in json using the following format. EXAMPLE START RESPONSE { "concept": "small vs big", "explanation": "The test image shows a small shape, similarly as all images on the left side. Conversely, the images on the right side feature big shapes.", "answer": "LEFT" } EXAMPLE END EXAMPLE START RESPONSE { "concept": "two circles vs one circle", "explanation": "The donut sits alone as the primary round element in the scene, matching the RIGHT examples where a single rounded object takes visual focus amid other details, like the soup bowl or drum.", "answer": "RIGHT" } EXAMPLE END RESPONSE
Prompt 7: Description-to-Side (D1S) task prompt. Use the descriptions below to solve the provided Bongard Problem. Respond in JSON using the following format. EXAMPLE START REQUEST { "left_descriptions": [ "A wired telephone with a coiled cord resting on a wooden table", "Power lines stretching across a clear blue sky with birds perched on them", "A clothes drying rack with colorful clothes hanging under the sunlight", "A swing hanging from a tree branch in a park with a child playing nearby", "A needle with thread passing through fabric, placed on a sewing table", "A pair of shoes with neatly tied laces placed on a tiled floor" ], "right_descriptions": [ "A smartphone with a cracked screen lying on a desk with visible damage", "A bird flying freely in the sky with its wings spread wide", "A pile of folded clothes neatly stacked on a shelf in a wardrobe", "A park bench surrounded by fallen leaves in an autumn setting", "A piece of fabric with intricate embroidery placed on a sewing table", "A pair of slip-on sandals placed on a carpeted floor near a doorway" ], "test_description": "A sailboat with its sails fully open gliding across a calm lake under a clear sky" } RESPONSE { "concept": "line is present vs no line is present", "explanation": "The first test image presents a sailboat with ropes hanging the sails. The ropes are lines, the same as items on the LEFT, like the telephone cord, power lines, and shoelaces.", "answer": "LEFT" } EXAMPLE END EXAMPLE START REQUEST { "left_descriptions": [ "Two bagels on a cutting board beside a knife", "A pair of headphones resting on a desk with the earcups facing up", "Two apples placed next to each other on a picnic blanket", "A pair of camera lenses lined up on a shelf", "A bicycle lying on the ground with both wheels visible", "A snowman with two stacked parts, one above the other" ], "right_descriptions": [ "Two forks placed on a plate", "A dartboard mounted on a brick wall with scores written on two papers beside it", "A spotlight shining onto a dark floor surrounded by cables", "A pie baking in the oven next to a tray of gingers bread men", "A record spinning on a turntable with a hand reaching toward it", "A teacup seen from above with steam rising", ], "test_description": "An octagonal dessert plate holding a single donut with sprinkles" } RESPONSE { "concept": "two circles vs one circle", "explanation": "The donut sits alone as the primary round element in the scene, matching the RIGHT examples where a single rounded object takes visual focus amid other details, like the soup bowl or drum.", "answer": "RIGHT" } EXAMPLE END REQUEST <request> RESPONSE
Prompt 8: Shared task introduction used in both I2S and D2S settings. You are a vision understanding module (an expert) designed to provide short, clear and accurate answers. The goal in solving a Bongard Problem is to classify each of the two test images to the corresponding class. There are two classes: LEFT and RIGHT. Each image belongs to exactly one class (either LEFT or RIGHT). All images belonging to the LEFT class represent a common, shared concept which is not present in any image from the RIGHT class, and vice versa - all images belonging to the RIGHT class represent a common, shared concept which is not present in any image from the LEFT class. The two test images belong to different classes.
Prompt 9: Images-to-Sides (I2S) task prompt. Solve the provided Bongard Problem. Respond in json using the following format. EXAMPLE START RESPONSE { "concept": "small vs big", "first": { "explanation": "The test image shows a small shape, similarly as all images on the left side. Conversely, the images on the right side feature big shapes.", "answer": "LEFT" }, "second": { "explanation": "The test image shows a big shape, similarly as all images on the right. The images on the left, on the other hand, feature small shapes.", "answer": "RIGHT" } } EXAMPLE END EXAMPLE START RESPONSE { "concept": "two circles vs one circle", "first": { "explanation": "The donut sits alone as the primary round element in the scene, matching the RIGHT examples where a single rounded object takes visual focus amid other details, like the soup bowl or drum.", "answer": "RIGHT" }, "second": { "explanation": "The tennis net and two balls represent two circular elements in the scene, aligning with the LEFT examples where pairs of circular objects are present, such as the bagels, apples, or camera lenses.", "answer": "LEFT" } } EXAMPLE END RESPONSE
Prompt 10: Descriptions-to-Sides (D2S) task prompt. (1/2) Use the image and the descriptions below to solve the provided Bongard Problem. Respond in JSON using the following format. EXAMPLE START REQUEST { "left_descriptions": [ "A wired telephone with a coiled cord resting on a wooden table", "Power lines stretching across a clear blue sky with birds perched on them", "A clothes drying rack with colorful clothes hanging under the sunlight", "A swing hanging from a tree branch in a park with a child playing nearby", "A needle with thread passing through fabric, placed on a sewing table", "A pair of shoes with neatly tied laces placed on a tiled floor" ], "right_descriptions": [ "A smartphone with a cracked screen lying on a desk with visible damage", "A bird flying freely in the sky with its wings spread wide", "A pile of folded clothes neatly stacked on a shelf in a wardrobe", "A park bench surrounded by fallen leaves in an autumn setting", "A piece of fabric with intricate embroidery placed on a sewing table", "A pair of slip-on sandals placed on a carpeted floor near a doorway" ], "first": "A sailboat with its sails fully open gliding across a calm lake under a clear sky", "second": "A launch boat with an engine cruising through the water leaving a trail of waves behind" } RESPONSE { "concept": "line is present vs no line is present", "first": { "explanation": "The first test image presents a sailboat with ropes hanging the sails. The ropes are lines, the same as items on the LEFT, like the telephone cord, power lines, and shoelaces.", "answer": "LEFT" }, "second": { "explanation": "The second image shows a motorized launch boat, which typically doesn’t have sails or ropes visible. This aligns more with the RIGHT side examples where no lines are present, like the smartphone, slip-on sandals, or park bench.", "answer": "RIGHT" } } EXAMPLE END
Prompt 11: Descriptions-to-Sides (D2S) task prompt. (2/2) EXAMPLE START REQUEST { "left_descriptions": [ "Two bagels on a cutting board beside a knife", "A pair of headphones resting on a desk with the earcups facing up", "Two apples placed next to each other on a picnic blanket", "A pair of camera lenses lined up on a shelf", "A bicycle lying on the ground with both wheels visible", "A snowman with two stacked parts, one above the other" ], "right_descriptions": [ "Two forks placed on a plate", "A dartboard mounted on a brick wall with scores written on two papers beside it", "A spotlight shining onto a dark floor surrounded by cables", "A pie baking in the oven next to a tray of gingers bread men", "A record spinning on a turntable with a hand reaching toward it", "A teacup seen from above with steam rising", ], "first": "An octagonal dessert plate holding a single donut with sprinkles", "second": "A tennis net with two balls lying nearby" } RESPONSE { "concept": "two circles vs one circle", "first": { "explanation": "The donut sits alone as the primary round element in the scene, matching the RIGHT examples where a single rounded object takes visual focus amid other details, like the soup bowl or drum.", "answer": "RIGHT" }, "second": { "explanation": "The tennis net and two balls represent two circular elements in the scene, aligning with the LEFT examples where pairs of circular objects are present, such as the bagels, apples, or camera lenses.", "answer": "LEFT" } } EXAMPLE END REQUEST <request> RESPONSE
Figure 13: Bongard-RWR+. (a) Left: Figures elongated vertically. Right: Figures elongated horizontally. (b) Left: Smooth contour figures. Right: Twisting contour figures.
Figure 14: Bongard-RWR+/GS. (a) Left: Extensions of segments cross at one point. Right: Extensions of segments do not cross at one point. (b) Left: Three parts. Right: Four parts.
Figure 15: Bongard-RWR+/L2. (a) Left: Empty. Right: Not empty. (b) Left: Large figures. Right: Small figures.
Figure 16: Bongard-RWR+/L3. (a) Left: Vertical hatched lines. Right: Horizontal hatched lines. (b) Left: Triangles. Right: Quadrangles.
Figure 17: Bongard-RWR+/L4. (a) Left: An acute angle directed inward. Right: No angle directed inward. (b) Left: Neck horizontal. Right: Neck vertical.
Figure 18: Bongard-RWR+/L5. (a) Left: One figure. Right: Two figures. (b) Left: More solid black circles. Right: More outline circles.
Figure 19: Bongard-RWR+/L6. (a) Left: A large hole. Right: A small hole. (b) Left: Shading thicker on the right side. Right: Shading thicker on the left side.